# suppSupplementary Material References

### v1

@@ Abstract
Autocrats use secret police to stay in power, as these organizations deter and suppress opposition to their rule. Existing research shows that secret police are very good at this but, surprisingly, also that they are not as ubiquitous in autocracies as one may assume, existing in less than 50% of autocratic country-years. We thus explore under which conditions secret police emerge in dictatorships. For this purpose, we apply statistical variable selection techniques to identify which of several candidate variables extracted from the literature on state security forces and authoritarian survival hold explanatory power. Our results highlight that secret police are more likely to emerge when rulers face specific, preempt-able threats, such as protests and anti-system mobilisation, but also when they have the material resources to establish these organisations. This research contributes to our understanding of autocrats’ institutional choices and authoritarian politics.
@@
Notes: We thank Janina Beiser-McGrath, Olga Vlasova, and the audience at the 2025 Political Studies Association Conference, Birmingham, for their comments.
@@ 1 Introduction
Secret police are a key instrument in the autocrat’s quest to suppress opposition and remain in power. They are specialists in surveillance and preventive repression, as they are tasked with instilling fear, deterring dissident political mobilisation, and, where such mobilisation takes place, putting an end to it before it can escalate to threaten the ruler. Accordingly, existing work has shown that secret police presence is associated with reductions in both individual and collective acts of resistance by citizens (Hager and Krakowski, 2022, 2024; Choulis et al., 2024). While they can be broadly linked to increased physical repression (Mehrl and Choulis, 2024), there also is evidence that secret police succeed at deterring dissent. For instance in Eastern Germany, their presence is linked to reduced levels of political imprisonment (Steinert, 2023), at least once they have established a reputation for tracking down and repressing opposition activity (Mehrl and Choulis, 2024).
@@
Recent research has thus begun to elucidate how secret police keep autocratic rulers in power. Earlier work has studied the principal-agent relationship between ruler and secret police, highlighting autocrats’ incentives to limit the competence of their subordinates (Egorov and Sonin, 2011; Zakharov, 2016; Dragu and Przeworski, 2019) but also the career pressures pushing subordinates to join the secret police in the first place (Scharpf and Gläßel, 2020, 2022). What may get lost across this body of work, however, is that secret police are not an automatic feature of autocratic governance. Indeed, global data shows that they exist in approximately a quarter of all military and personalist regime-years coded by Geddes, Wright, and Frantz (2014). They are present in less than half of all party regime-years, have been highly prevalent only in the dictatorships of Europe and Central Asia, and, overall, existed in not even a third of all non-democracies over the period 1950–2018 (Choulis et al., 2024; Mehrl and Choulis, 2024). Given that secret police apparently generally succeed in preventing dissent, this raises the question – when can we expect secret police to be formed in autocracies?
@@
Here, we answer this question. Drawing on existing research on states’ security force design, we first identify a number of potential predictors of secret police existence. We then apply a LASSO model (Tibshirani, 1996) to select the subset of these predictors that are associated with the onset of secret police organisations. This selection occurs by penalizing the number of non-zero entries in the maximized likelihood. Coefficients with little impact on the target variable are set to zero, effectively removing those predictors. This allows us to identify the variables that have the strongest influence on the formation of secret police. We find that secret police are most likely to be established when rulers not only perceive specific, preempt-able rather than acute, threats that drive their creation, but also possess the necessary material resources to act upon them.
@@
This research advances the literature on authoritarian security institutions and autocracy more generally in several ways. First, we collect different theoretical arguments regarding the drivers of security institutional set-ups from existing studies and test them against each other in a principled fashion, finding that some prominent explanations contribute little to understanding secret police formation. Second, we provide the first study of when rulers choose to institute secret police organisations, thereby, third, providing a clear foundation for further research that may investigate the specific role of single predictors of secret police from a more theoretical and/or causal perspective. Finally, our research emphasizes that what has regularly been considered a static feature of autocratic governance – secret police – is actually a variable to be explained, thus highlighting important variation in authoritarian institutional choices.
@@ 2 The Predictors of State Security Force Structure
Existing research has identified several predictors of state security force structure, that is, how governments decide to structure their security apparatus and what types of institutions they choose to invest in. This research has, in particular, studied the existence of counterbalancing paramilitaries and pro-government militias, but has also begun to investigate when government recruit private military contractors and foreign legionnaires. And while these types of security forces may share commonalities, but also important differences, with secret police organisations, their identified predictors provide a useful starting point to understand when rulers choose to establish a secret police.
Perhaps unsurprisingly, a first set of drivers of state security force structure very directly pertain to the risk of deposition the ruler is facing. Along these lines, several studies find that facing a higher risk of being deposed by regime elites via a coup d’etat, i.e., the level of internal threat, increases rulers’ extent of coup-proofing, via the establishment of paramilitaries such as presidential guards (Belkin and Schofer, 2003, 2005) or pro-government militias (Ash, 2016; Carey et al., 2016), but also by hiring private military contractors (Gentil-Fernandes et al., 2024). But beyond coup risk, research also argues that rulers are susceptible to threats which are external to their regime, in particular mass-based mobilisation challenges in the form of protests or intrastate conflict (Carey et al., 2016; Böhmelt and Clayton, 2018; Akins, 2021; Klosek and Souleimanov, 2025), or even their country, pointing to the role of interstate rivalries, militarized interstate disputes (MIDs), and wars (Ash, 2016; Akins, 2021; Grasmeder, 2021). And Böhmelt et al. (2017) focus again on the role of coup risk, but show that regimes also learn external events as they react to the experience and behaviour of their “peers”, that is, countries facing similar threats.
@@
Focusing more on the constraints rulers face, Böhmelt and Clayton (2018) highlight that specific security force configurations require more developed state capacity on the side of governments, arguing in particular that mobilising, maintaining, and supervising paramilitaries requires more financial and administrative resources than linking up with militias would entail. Mehrl and Choulis (2021), in turn, emphasise that employing paramilitaries was more straightforward for countries inheriting similar auxiliary forces upon colonial independence, while research on militias and private military contractors points to the benefits of delegating violence to such “external” actors when accountable to democratic aid donors (Carey et al., 2015, 2016; Gentil-Fernandes et al., 2024).
@@ 3 The (Potential) Predictors of Secret Police
As we aim to understand when rulers establish secret police organisations, the insights generated by the literature on state security force structure guide the choice of variables included in our models. Therefore, we include several indicators of events directly threatening the ruler, both “at home” but also in their neighborhood, regime type, and financial and administrative state capacity (see Table 1). However, we do not account for colonial history and donor accountability. The reason for this is that, first, secret police are unlikely to be inherited across regimes (Geddes et al., 2018; Mehrl and Choulis, 2024). And second, secret police, as official state institutions whose effect depends on citizens actually knowing about them, should offer little plausible deniability.
Instead, we add a number of covariates that capture societal risk factors associated with anti-government mobilisation. These attributes should increase the ruler’s risk perception and, accordingly, willingness to invest into a secret police. Therefore, we incorporate covariates regarding the population size, the population share of politically included and excluded ethnic groups, as well as what percentage of the population lives in urban settings. These variables have been prominently linked to regime threats and survival strategies (see e.g. Roessler, 2011; Chenoweth and Ulfelder, 2017). Similarly, variables counting the time the current leader and regime have been in power are included as both the risk of deposition and the associated need to set up new security measures are likely to change across their life cycles as the ruler becomes more established (Svolik, 2012). We also include a measure capturing whether a regime uses ideological legitimization claims, following the notion that such regimes may have reduced needs for security institutions (Escribà-Folch et al., 2020) or, alternatively, feel a stronger need to protect their core ideological tenets among their citizens.
We also seek to capture more formal precursors of political mobilisation. Therefore, we include a dummy variable for election years as well as measures of civil society independence, participation level, and the presence of anti-government civil society organisations. We account for the ruler’s existing efforts at tackling such threats by including covariates that gauge the level of repression, both generally and targeted specifically at civil society organisations. Given that rulers only have a limited amount of resources to create and sustain different security institutions (Böhmelt and Clayton, 2018), variables capturing alternative security force structures, clientelistic rule as an alternative mode of generating information and control, as well as overall military spending are included as possible predictors. Expanding on the intuition that rulers require resources, i.e., disposable income, to create and maintain security forces (Geddes et al., 2018, p.157; Fails, 2020), we include measures of economic performance and growth, and of oil and gas production.
@@
Table 1 lists all covariates included in our models [^1] Table 1 also gives data sources. In accessing data, we benefited from Peacesciencer (Miller, 2022) and the WEP Dataverse (Graham and Tucker, 2019).. Our goal here is to provide a first indication of which of these factors, and, thus, the theoretical dynamics discussed above, are actually associated with the establishment of secret police institutions. We thus do not develop a specific, fully formulated theory for each of them, but instead lay the groundwork for such theoretical development by investigating which of the several potential theoretical dynamics actually appear worthy of investigation [^2] Additionally, we opt against developing theoretical hypotheses here as testing all of them together would likely prove close to impossible (see Keele et al., 2020)..
@@
Intrastate conflict: Dummy (UCDP)1 - Clientelism (V-Dem)6 -
Intrastate conflict: Years since (UCDP)1 - Ethnically excluded population (%)10 -
Democracy score (Polity5)2 - Counterbalancing11 -
Coup attempt: Dummy3 - Military expenditures (Latent score, ln)12 1, 2, 3, 4
Coup attempt: Years since3 - Oil production (Financial value, ln)13 -
Human rights (Latent score)4 1, 2 Gas production (Financial value, ln)13 3, 4
State capacity (Latent score)5 - Election year (NELDA)14 -
Bureaucratic capacity (V-DEM)6 - Leader duration (CHISOLS)7 -
Fiscal capacity (V-DEM)6 - Regime duration (CHISOLS)7 1, 2, 3, 4
Territorial control (V-DEM)6 - Failed coups in region3 -
CSO entry and exit (V-DEM)6 - Successful coups in region3 -
CSO repression (V-DEM)6 1, 2, 3, 4 International rivalry: Dummy15 1, 2, 3, 4
CSO participatory environment (V-DEM)6 1, 2 International rivalry: Count15 1, 2, 3, 4
CSO anti-system movements (V-DEM)6 1, 2, 3, 4 MID: Dummy16 -
CSO strength (V-DEM)6 - MID: Years since16 -
Regime change (CHISOLS)7 - MID: Count16 3, 4
Urban population (%)8 - Protest (Latent score)17 1, 2, 3, 4
Economic growth8 - Neighbour protest (Latent score)17 -
Personalisation (Latent score)9 3, 4 GDP p.c. (Logged)8 1, 2, 3, 4
Ideological legitimisation (V-Dem)6 3, 4 Population (Logged)8 -
  • Data sources: 1Davies et al. (2024); 2Marshall and Gurr (2020); 3Powell and Thyne (2011); 4Fariss (2019); 5Hanson and Sigman (2021); 6Coppedge et al. (2023); 7Mattes et al. (2016); 8World Bank (2021); 9Geddes et al. (2018); 10Vogt et al. (2015); 11Pilster and Böhmelt (2011); 12Barnum et al. (2025); 13Ross and Mahdavi (2015); 14Hyde and Marinov (2012); 15Thompson and Dreyer (2011); 16Palmer et al. (2022); 17Chenoweth et al. (2014).

### v2

# Guardians of the Regime: Secret Police Formation in Autocracies

@@ Abstract
Autocrats use secret police to stay in power, as these organizations deter and suppress opposition to their rule. Existing research shows that secret police succeed at this but, surprisingly, also that they are not as ubiquitous in autocracies as one may assume, existing in fewer than half of autocratic country-years. We thus explore under which conditions secret police emerge in dictatorships. For this purpose, we develop a theoretical framework for potential predictors and apply statistical variable selection techniques to identify which of several candidate variables extracted from the literature on state security forces and authoritarian survival hold explanatory power. Our results highlight that secret police are more likely to emerge when rulers face structural, regime-external threats, such as organised anti-system mobilisation and international rivals, or witness successful regime-internal contestation abroad that hints at similar threats at home. But additionally, we find that rulers must have sufficient material resources and personalised power to establish secret police. This research contributes to our understanding of autocrats’ institutional choices and authoritarian politics.
@@
Notes: We thank Janina Beiser-McGrath, Olga Vlasova, the audience at the 2025 Political Studies Association Conference, Birmingham, and the anonymous reviewers and editor for their comments.
@@ 1 Introduction
Secret police are a key instrument in the autocrat’s quest to suppress opposition and remain in power. They are specialists in surveillance and preventive repression, as they are tasked with instilling fear, deterring dissident political mobilisation, and, where such mobilisation takes place, putting an end to it before it can escalate to threaten the ruler. Accordingly, existing work has shown that secret police presence is associated with reductions in citizens’ individual and collective resistance (Hager and Krakowski, 2022, 2024; Choulis et al., 2024). While they can be broadly linked to increased physical repression (Mehrl and Choulis, 2024), there also is evidence that secret police succeed at deterring dissent. For instance in Eastern Germany, their presence is linked to reduced levels of political imprisonment (Steinert, 2023), at least once they have established a reputation for tracking down and repressing opposition activity (Mehrl and Choulis, 2024).
Recent research has thus begun to elucidate how secret police, such as the Stasi, Pinochet’s DINA/CNI, and Assad’s Political Security Directorate, keep autocratic rulers in power. Secret police can be defined as “an internal security agency that uses violent policing practices and intelligence operations against political opponents or dissidents while being autonomous from the rest of the security apparatus” (Mehrl and Choulis, 2024, p.995). They thus differ from internal security institutions that report to the military, conduct operations combining violent policing and intelligence against criminal rather than political targets, target political opponents exclusively through intelligence work, or, like paramilitaries and militias, combine violence against such opponents with minimal intelligence activity [^1] Additionally, militias differ from secret police due to their informal links to the state (Carey et al., 2015). (see Mehrl and Choulis, 2024; Plate and Darvi, 1982).
@@
Earlier work has studied the principal-agent relationship between ruler and secret police, highlighting autocrats’ incentives to limit the competence of their subordinates (Egorov and Sonin, 2011; Zakharov, 2016; Dragu and Przeworski, 2019; Thomson, 2020), but also the career pressures pushing subordinates to join secret police such as Argentina’s Battalion 601 in the first place (Scharpf and Gläßel, 2020, 2022). More recently, research investigates how secret police navigate the trade-off between using the intelligence supplied by informants and the risk of their exposure (Liu et al., 2026). What may get lost across this body of work, however, is that secret police are not an automatic feature of autocratic governance. Indeed, global data shows that they exist in approximately a quarter of all military and personalist regime-years coded by Geddes et al. (2014). They are present in less than half of all party regime-years, have been highly prevalent only in the dictatorships of Europe and Central Asia, and, overall, existed in less than a third of all non-democracies over the period 1950–2018 (Choulis et al., 2024; Mehrl and Choulis, 2024). Given that secret police generally succeed in preventing dissent, this raises the question – when and where can we expect secret police to be formed in autocracies?
@@
Here, we answer this question. We develop a framework that distinguishes potential predictors of secret police formation along three dimensions – willingness–opportunity, structure–shock, and regime internal–regime external – and then identify candidate predictors from existing research on states’ security force design and authoritarian survival. We apply the least absolute shrinkage and selection operator (LASSO, Tibshirani, 1996) to select the subset of these predictors that are associated with the onset of secret police organisations. This selection occurs by penalising the sum of the absolute values of the coefficients often leading to a number of non-zero entries. Coefficients with little impact on the target variable are set to zero, effectively removing those predictors. This allows us to identify the variables that have the strongest influence on the formation of secret police. Here, we apply a modified Lasso approach where stratification according to the outcomes in the cross-validation step is employed to ensure that the model is stable for binary outcomes with rare events, such as secret police formation. We find that secret police are most likely to be established when rulers not only perceive specific, structural rather than acute threats from outside of the regime, but also possess the necessary material resources and power base to act upon these threats.
@@
This research advances the literature on authoritarian security institutions and autocracy more generally in several ways. First, we collect different theoretical arguments regarding the drivers of security institutional set-ups from existing studies, develop a theoretical framework to categorise them, and test them against each other in a principled fashion, finding that some prominent explanations contribute little to understanding secret police formation. Second, we provide the first study of when rulers choose to institute secret police organisations, thereby, third, providing a clear foundation for further research that may investigate the specific role of single predictors of secret police from a more theoretical and/or causal perspective. Finally, our research emphasises that what has regularly been considered a static feature of autocratic governance – secret police – is actually a variable to be explained, thus highlighting important variation in authoritarian institutional choices.
@@ 2 The Predictors of State Security Force Structure
Existing research has identified several predictors of state security force structure, that is, how governments decide to structure their security apparatus and what types of institutions they choose to invest in. This research has, in particular, studied the existence of counterbalancing paramilitaries and pro-government militias, but has also begun to investigate when governments recruit private military contractors and foreign legionnaires. As these types of security forces may share commonalities with secret police, but also important differences such as their main targets, emphasis on intelligence versus on violence, and clarity of links to the regime, their identified predictors provide a useful starting point to understand when rulers choose to establish a secret police.
Perhaps unsurprisingly, a first set of drivers of state security force structure very directly pertain to the risk of deposition the ruler is facing. Along these lines, several studies find that facing a higher risk of being deposed by regime elites via a coup d’etat, i.e., the level of internal threat, increases rulers’ extent of coup-proofing, via the establishment of paramilitaries such as presidential guards (Belkin and Schofer, 2003, 2005) or pro-government militias (Ash, 2016; Carey et al., 2016), but also by hiring private military contractors (Gentil-Fernandes et al., 2024). But beyond coup risk, research also argues that rulers are susceptible to threats which are external to their regime, in particular mass-based mobilisation challenges in the form of protests or intrastate conflict (Carey et al., 2016; Böhmelt and Clayton, 2018; Akins, 2021; Klosek and Souleimanov, 2025), or even their country, pointing to the role of interstate rivalries and conflict (Ash, 2016; Akins, 2021; Grasmeder, 2021). And Böhmelt et al. (2017) focus again on the role of coup risk, but show that regimes also learn from external events as they react to the experience and behaviour of their “peers”, that is, countries facing similar threats.
@@
Finally, focusing on the constraints rulers face, Böhmelt and Clayton (2018) highlight that governments require more developed state capacity for specific security force configurations, arguing that mobilising, maintaining, and supervising paramilitaries requires more financial and administrative resources than linking up with militias would entail [^2] Focusing on where states get their financial resources from, research on militias and private military contractors argues that reliance on democratic aid donors makes delegating violence to such “external” actors attractive (Carey et al., 2015, 2016; Gentil-Fernandes et al., 2024). However, secret police offer no such plausible deniability. Similarly, colonial inheritance is one source of paramilitaries, but secret police generally do not transfer from one regime to another (Mehrl and Choulis, 2021, 2024). We therefore do not consider these potential predictors below..
@@ 3 The (Potential) Predictors of Secret Police
The insights generated by the literature on state security forces offer a starting point to identify factors that push rulers towards or away from establishing secret police. To facilitate this identification and the eventual systematic theorisation of factors, we distinguish them along three dimensions. First, willingness and opportunity (Most and Starr, 1984): is the variable expected to affect the ruler’s motivation to establish a secret police, the constraints the ruler may have on this policy choice, or both? Second, structure versus disruptive events (Bermeo, 2016; Timoneda et al., 2023): does a variable capture structural conditions which may affect the baseline probability of secret police formation, and gradually shift it over time, or a shock-like event indicating instantaneous changes in the dictator’s willingness or opportunity for secret police formation? Third, regime versus external drivers (Bueno de Mesquita et al., 2003; Roessler, 2011): is a factor located within the regime itself, for instance as a result of elite bargaining, power distribution, or regime emergence, or does it relate to the attributes and actions of regime outsiders?
These three dimensions offer a framework to think about, identify, and categorise the potential predictors of secret police formation, and more generally of policy shifts within the arena of authoritarian politics. Accordingly, the distinction between willingness and opportunity reflects what payoffs a leader expects, and what constraints they face, when forming a secret police. This phenomenon mirrors the distinction between militaries’ disposition and ability to carry out a coup d’etat (Finer, 1962; Powell, 2012). Here, willingness captures that secret police may assist dictators in navigating a challenging threat environment (Hager and Krakowski, 2022; Choulis et al., 2024), but also impose political costs and present threats to the dictator (Egorov and Sonin, 2011; Thomson, 2020). The opportunity dimension captures that dictators, even if willing to form a secret police, may be constrained to do so by their position vis-a-vis other elites (Svolik, 2012) but also a lack of material resources to train and maintain costly security forces (Böhmelt and Clayton, 2018). As such, the corresponding first dimension covers several dynamics ascribed to authoritarian decision-making, and captures whether a potential predictor of secret police affects the rulers’ motivation or obstacles to form such an institution.
In line with recent research on authoritarian politics (Timoneda et al., 2023), the second and third dimension indicate whether it is a structural feature of or more of a shock-like shift to a given autocratic context, and whether it pertains to regime insiders or to actors and dynamics that are outside of it. This framework, as illustrated in Table 1 with some of the variable discussed below, enables the identification and selection of variables to be included in the empirical models in a structured fashion, and facilitates model interpretation. Additionally, it provides a bridge point for the study of potential further predictors of secret police emergence, allowing for them to be straightforwardly connected to the insights of the models we present here.
Table 1: (Potential) predictors of Secret Police along three dimensions.
Structure Shock
Willingness Regime & leader duration Coup attempt
Opportunity Personalism; State capacity Coup attempt
(a) Regime-internal predictors
Structure Shock
Willingness Excluded ethnic population; substitute security forces Protests; Neighbour coups & protests
Opportunity Wealth; Oil & gas production Economic Growth
(b) Regime-external predictors
@@
That being said, our goal here is to provide a first indication of which of these factors, and, thus, the theoretical dynamics discussed below, are associated with the establishment of secret police institutions. We thus do not develop a specific, fully formulated theory for each of them. Instead, we lay the groundwork for such theory development via the framework introduced here and exploring which of the several potential theoretical dynamics actually appear worthy of investigation [^3] We also opt against developing theoretical hypotheses as testing them together would likely prove impossible (Keele et al., 2020)..
In line with earlier research on security force structure, we include several indicators of events directly threatening the ruler in our model. For instance, coup attempts are regime-internal shocks that may shift the ruler’s willingness and opportunity to form a secret police, while other threats at home, such as protests, civil and interstate conflict, or those occurring in their neighbourhood represent external shocks. In contrast, autocratic regime type and financial and administrative state capacity represent structural factors linked to the regime which should mainly affect the dictator’s opportunity to institute secret police, but leave their willingness to do so comparatively unaltered. These factors are placed accordingly into Table 1 and we list their operationalization into variables to be included in our models in Table 2.
Beyond these potential predictors taken from existing research on state security forces, we next turn to the broader literature on authoritarian politics and survival to identify further potential factors within and outside of the regime, structural or shock-like, that may affect dictator’s opportunity and/or willingness to create secret police. First, we include several structural covariates that capture societal risk factors associated with anti-government mobilisation, which should increase the ruler’s risk perception and, accordingly, willingness to invest into a secret police. Therefore, we incorporate covariates regarding the population size, the population share of politically included and excluded ethnic groups, as well as what percentage of the population lives in urban settings. These variables have been prominently linked to regime threats and survival strategies (see e.g. Roessler, 2011; Chenoweth and Ulfelder, 2017). Along the same lines, but capturing closer structural precursors of political mobilisation, we also include measures of civil society independence, participation level, and the presence of anti-government civil society organisations. And as rulers may well expect mass mobilisation, and thus the need for secret police, to increase disproportionately around elections, we include a dummy variable for election years.
Similarly, we add further structural characteristics of the regime that capture how established and cohesive the dictator’s rule is, and hence should affect their perceived need and according willingness to create secret police. Along these lines, we include variables counting the time the current leader and regime have been in power (Svolik, 2012). We include a measure capturing whether a regime uses ideological legitimisation claims, following the notion that such regimes may have reduced needs for security institutions as their stability arises from shared goals (Escribà-Folch et al., 2020) [^4] Note that alternatively, ideological legitimisation claims may also trigger a stronger need to protect their core ideological tenets among their citizens, and thus increase regime willingness to form secret police.. And given that the ruler’s expected payoffs from, and thus willingness to invest in, another repressive agent likely depend on their existing efforts at tackling threats, we include covariates that gauge the level of repression, both generally and targeted specifically at civil society organisations.
In light of this, we also include as a structural regime feature the existence of other security institutions with relatively similar tasks as secret police – counterweight paramilitaries, affiliated forces, and the overall degree of counterbalancing. Given that all of these actors can, at least to some degree, be used to counter threats emanating from within the regime as well as other domestic sources of danger, they may substitute for each other, with the presence of one reducing the ruler’s need to create another. For the same reason, we include a variable capturing clientelistic rule, which offers an alternative mode of generating information and control, thus reducing the need for a formal secret police.
In theory, a ruler may also choose to create endless parallel security institutions to counterbalance each other, but forming even one such institution requires extensive resources (Böhmelt and Clayton, 2018; Geddes et al., 2018, p.157). As such, we include several structural features of the regime’s resource environment, including the country’s general economic performance, defence spending, as well as oil and gas production as a source of disposable income (Fails, 2020). These structural characteristics should affect dictator’s opportunities to fund the formation of new secret police institutions, as should shock-like shifts in economic growth.
Table 2 lists all covariates included in our models, specifying how we operationalise the variables discussed above. Note that in several cases, we include various operationalizations and leave selecting among them to the model. We report data sources for all included variables in the supplementary materials.
@@
Intrastate conflict: Dummy - Clientelism -
Intrastate conflict: Years since - Ethnically excluded population (%) -
Neighborhood intrastate conflicts (count) - Counterweight Forces -
Democracy score - Counterbalancing -
Coup attempt: Dummy - Military expenditures (Latent score, ln) 1, 2, 3, 4
Coup attempt: Years since - Oil production (Financial value, ln) -
Human rights (Latent score) 1, 2 Gas production (Financial value, ln) 3, 4
State capacity (Latent score) - Election year -
Bureaucratic capacity - Leader duration 1, 2
Fiscal capacity - Regime duration 1, 2, 3, 4
Territorial control - Neighborhood unsuccessful coups (count) -
CSO entry and exit - Neighborhood successful coups (count) 1, 2, 3, 4
CSO repression 1, 2, 3, 4 International rivalry: Dummy 3, 4
CSO participatory environment 1, 2 International rivalry: Count 1, 2
CSO anti-system movements 1, 2, 3, 4 MID: Dummy -
CSO strength - MID: Years since -
Regime change - MID: Count 1, 2, 3, 4
Urban population (%) - Protest (Latent score) 1, 2, 3, 4
Economic growth - Neighborhood protest (average score) 1, 2
Personalisation (Latent score) 1, 2, 3, 4 GDP p.c. (ln) 1, 2, 3, 4
Ideological legitimisation 3, 4 Population (ln) -
Affiliated forces - Neighborhood: Secret Police (count) -